3. A ladder 5.10 m long leans against a wall inside a spaceship. From the point of view of a person on the ship, the top of the ladder is 4.65 m above the floor. The spaceship moves past the Earth with a speed of 0.82c in a direction parallel to the floor of the ship. What is the length of the ladder as seen by an observer on Earth? m
Answer #1

given

ladder 5.10 m long leans against a wall inside a spaceship

ladder is 4.65 m above the floor

A2 = B2 + C2

5.12 = 4.652 + C2

C = 2.095 m

L ' = L ( 1- (v/C)2 )1/2

= 2.095 x ( 1 - (0.82C/C)2 )1/2

= 2.095 x ( 1 - (0.82)2 )1/2

so the length of the ladder as seen by an observer on Earth is

L ' = 1.1991 m